if 10x+5=6 then x=?

Grade:12

30 Answers

OUTPASS
8 years ago
1/10=x
RAJU
8 years ago
1/10
charan
8 years ago
1/10
Nandini Menon
8 years ago
10x+5=6
10x=6-5
x=1-10
therefore, x=-9
pa1
8 years ago
 
 
10x+5=6
10x=6-5
x=1-10
therefore, x=-9
ROHIT
8 years ago
1/10=0.1
Mondo Camry
8 years ago
10x=6-5,
10x=1,
x=1/10,
x=0.1
 
 
plz approve
sri tanish
8 years ago
0.1
SHANMUKESHWAR
8 years ago
1/10
4222
8 years ago
1/10
SEELA YUVARAJU
8 years ago
1/10
sushil
6 years ago
tto get the answer follow these steps,
  1. ​10x+5=6
  2. take 5 to the other side, 10x=6-5
  3. 10x=1
  4. take 10 to the othe side,x=1/10
  5. the answer is 1/10 or0.1
yathartha gupta
6 years ago
The given equation is10x+5 = 6................110x = 6-510x = 1X = 1/10 So the x stands for 1/10 in given equation.
Harshit Panwar
6 years ago
10x+5=6
10x=6-5
10x=1
x=1/10
x=0.1
Therefore 1/10 or 0.1 is the answer for this question.
Hope it helps u.
PRANAV KUMAR
6 years ago
10x+5=6
10x=6-5
10x=1
x=1/10
x=0.1
 
ans=x=1/10 or 0.1
 
Hari Sundar Pyda
6 years ago
10x+5=6Bring 5 to the right hand side and change the sign10x= 6-510x=1Bring the 10 to the other side in the denominator. x=1/10Hope it helps!!
Subhash
6 years ago
in the given equation--
10X + 5 =6
therefore, 10X = 6-5
therefore, 10x=1
therefore,  x= 1/10
or x= .1
thats all
 
Kochouseph
6 years ago
10x+5=6 10x=6-5=1 x=1/10=0.1 This is the final answer. I hope it will help you. This is the simplest way.
Samarth Akula
6 years ago
10 x + 5 = 6 , so ,
10 x = 6 – 5,
10 x =1,
x = 1/10= 0.1
this is the simplefied way and i hope this will help you
Saket singh
6 years ago
To get the ans follow these steps..................10x+5=6so,Bring the 5 to the RHS ,it becomes negativeTherefore , 10x=6-5.....................10x=1........ .............x=1/10Hope u understand this
Amit
6 years ago
10x=6-5,10x=1,x=1/10,
Priyansh
6 years ago
10x +5 =6
 Transposing 5 to the right side of equation and changing sign we get
 
10x=6-5=1
Therefore 10x=1
Or x=1/10
       =0.1
Pradeep Tangirala
6 years ago
 The given equation is...10x + 5 = 6
So....10x= 6-5
                 =1
So...x=1/10 
      =0.1.... 
This is the only way to answer this question...
Mini James
6 years ago
10x + 5 = 6(Always keep the variable in one side of the equation and take all the other constants to the other side)First take +5 to the other side then the sign reverses i,e.10x =6 - 510x = 1X = 1/10
Om
5 years ago
Given ;
So  10x+5=6
So  10x=6-5=1        [as if we move a constant to another side of the equation its sign changes]
 So  10x=1
So  x=(1/10)             [as if we move a constant to another side of the equation its sign changes]
 
Therefore if 10x+5=6  then x would be (1/10).
Sai Soumya
5 years ago
dear student,
according to question , 10x+5=6
=>10x=6-5
=>10x=1
=>x=1/10
so the required answer is 1/10 or 0.1
hope this helps you.
Mustafa
5 years ago
Its question related to equation, according to me we will solve it like this....
Sai Soumya
5 years ago
according to the given question
10x=6-5
=>10x=1
=>x=1/10
So start by taking the constants one side and the variables the other side.
 
hope this helps you.
have a great day.
 
 
Atharv lomte
5 years ago
10x plus 5=6 then x=?
10x=6-5..........(by transposing 5 to r.h.s.)
10x=1
10x/10=1/10...(by dividing both sides by 10)
x=1/10
So our answer to this question is x=1/10
Sayantan Garai
5 years ago
10x+5=6
=>10x=6-5
=>10x=1
=>x=1/10
=>x=0.1(Ans.)
------------------------------------------------------------